OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 16 = 8x A. 2 B. 1 C. 0 I chose A..

OpenStudy (anonymous):

x^2-8x+16=(x-4)^2=0 x-4=0 x=4

sam (.sam.):

x^(2)-8x+16=0 Factor the trinomial x^(2)-8x+16. (x-4)(x-4)=0 Set each of the factors of the left-hand side of the equation equal to 0. x-4=0 Since -4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. x=4 Number of solutions: 1 Answer: B